SPISAK RADOVA Dr OLGE ATANACKOVIĆ- VUKMANOVIĆ

1. G. Teleki, O. Atanacković: The growth of knowledge on astronomical refraction I. 1983, Rep. No. 19, Dep. of Geodesy, Univ. of Uppsala, Sweden, 255-271.

2. Dj. Teleki, O. Atanacković: Rast znanja o astronomskoj refrakciji.

1984, Publ. astron. društva ”Rudjer Bošković ”, No.4, 151-155.

3. O. Atanacković: Preliminary investigation of the image motion at the Belgrade Observatory. 1985, Publ. Obs. Astron. Belgrade No.33, 83-87.

4. O. Atanacković, E. Simonneau: Effects of elastic collisions on the local frequency redistribution in the transfer of resonant line photons. 1985, Colloque ”Collisions et rayonnement”, Orleans, P3.

5. O. Atanacković: Le troisieme volume d’Epitome fundamentorum astronomiae. 1985, Bull. Inform. CDS No.28, 15-16.

6. O.Atanacković, E.Simonneau: Effects of elastic collisions on the local frequency redistribution in the transfer of resonant line photons. 1986, Annales de Physique, (Colloque No.3), Vol.11, 137-138.

7. O.Atanacković:Analiza nelokalnih efekata u ne-LTR prenosu zračenja u rezonantnim linijama zvezdanih atmosfera. 1986, Magistarski rad. Prirodno-matematički fakultet. Beograd.

8. O.Atanacković-Vukmanović, E.Simonneau: Kinetic effects in non-LTE line transfer in stellar atmospheric conditions. 1987, Bull. Obs. Astron. Belgrade 137, 66-79.

9. O.Atanacković-Vukmanović, E.Simonneau: Parameters characterizing non-LTE line radiative transfer in some astrophysical conditions. 1987, Bull. Obs. Astron. Belgrade 137, 58-65.

10. O.Atanacković, J.Borsenberger, J.Oxenius, E.Simonneau: Resonance line transfer and transport of excited atoms - III. Self-consistent solutions (2). 1987, J. Quant. Spectrosc. Radiat. Transfer 38, 427-446.

11. G.Teleki, O.Atanacković:On the history of astronomical refraction study. 1987, in ”Istoriko-astronomicheskie issledovaniya” , ed. A.A. Gurshtein, Nauka, Moskva, 139-150. (in russian).

12. O.Atanacković-Vukmanović, M.S.Dimitrijević, E.Simonneau: Karakteristike prenosa zračenja u pražnjenjima visokog pritiska koja se koriste u svetlosnim izvorima. 1987, Zbornik radova sa 29. Simpozijuma ”Etan u pomorstvu”, Zadar, 282-284.

13. O.Atanacković-Vukmanović, E.Simonneau: An approximative solution in the frame of kinetic non-LTE approach of Lyman alpha line transfer in chromospheric conditions. 1987, II Workshop ”Astrophysics in Yugoslavia”, Beograd, 21-22.

14. E.Simonneau, O.Atanacković: Iteracioni faktori u rešavanju ne-LTR prenosa zračenja u spektralnim linijama. 1988, IX nacionalna konferencija astronoma Jugoslavije, Sarajevo, (abstract).

15. O.Atanacković:Non-LTE radiative transfer.

1989, Bull. Obs. Astron. Belgrade 140, 127-130.

16. M.S.Dimitrijević, O.Atanacković-Vukmanović:Plasma screening effects on Stark broadening of ion lines at the adiabatic limit. 1989, Colloque ”Collisions et rayonnement”, Orleans.

17. B.Ševarlić, G.Teleki, O.Atanacković-Vukmanović, I.Pakvor: Bibliography of catalogues and papers on parallaxes. 1990, Publ. Obs. Astron. Belgrade 40, 79-128.

18. B.Ševarlić, G.Teleki, I.Pakvor, O.Atanacković-Vukmanović: Bibliography of catalogues and papers on proper motions. 1990, Publ. Obs. Astron. Belgrade 40, 129-238.

19. B.Ševarlić, G.Teleki, O.Atanacković-Vukmanović, I.Pakvor: Bibliography of catalogues and papers on radial velocities. 1990, Publ. Obs. Astron. Belgrade 40, 239-285.

20. M.S.Dimitrijević, O.Atanacković-Vukmanović:Plasma screening effects on Stark broadening of ion lines at the adiabatic limit. 1990, Annales de Physique, (Colloque No.3), Vol.15, 75-76.

21. E.Simonneau, O.Atanacković-Vukmanović: Iteration factors in the solution of the NLTE line transfer problem. 1991, in ”Stellar atmospheres: Beyond classical models”, eds. L. Crivellari et al., Kluwer Academic Publishers, 105–110.

22. O.Atanacković –Vukmanović : Prilog rešavanju problema prenosa zračenja u zvezdanim atmosferama 1991, Doktorska disertacija, Matematički fakultet, Univerzitet u Beogradu.

23. O.Atanacković –Vukmanović , E.Simonneau: Iteration factors method in the solution of the NLTE line transfer problem 1991, Proc. of the 1st General Conf. of the Balkan Physical Union, 138–140.

24. O.Atanacković –Vukmanović , E.Simonneau: The use of iteration factors in the line formation problem with spatial variations in profile function 1993, Publ. Obs. Astron. Belgrade 44, 41–44.

25. O.Atanacković – Vukmanović , E.Simonneau: The use of iteration factors in the solution of the NLTE line transfer problem – I. Two–level atom 1994, J. Quant. Spectrosc. Radiat. Transfer Vol. 51, No.3, 525–543.

26. L.Č. Popović, I.Vince, A. Kubičela, S. Samurović, O.Atanacković –Vukmanović: Contribution of gravitational redshift to spectral line profiles of Seyfert galaxies and quasars: The case of Lorentz profile 1994, Bull. Astron. Belgrade 149, 9–14.

27. O.Atanacković –Vukmanović , L.Č. Popović, I.Vince, A. Kubičela: Contribution of gravitational redshift to spectral line profiles of Seyfert galaxies and quasars: The case of Voigt profile 1994, Bull. Astron. Belgrade 150, 1–5.

28. L.Č. Popović, I. Vince, O.Atanacković–Vukmanović, A. Kubičela: Contribution of gravitational redshift to spectral line profiles of Seyfert galaxies and quasars 1995, Astron. Astrophys. 293, 309–314.

29. O.Atanacković–Vukmanović, E. Simonneau: On some iteration factors families in solution of line transfer problem with depth–dependent profile function 1995, Publ. Obs. Atron. Belgrade 49, 77–80.

30. I. Vince, L. Č. Popović, S. Jankov, G. Djurašević, O. Atanacković –Vukmanović , D. Jevremović : On inverse methods used at Belgrade Observatory for analysis of spectral line shapes 1995, Publ. Obs. Astron. Belgrade 50, 21–26.

31. L. Č. Popović, I. Vince, S. Jankov, G. Djurašević, O. Atanacković –Vukmanović , D. Jevremović : Analysis of the MgII h spectral line shapes in HR7275 and IM Peg 1995, Publ. Obs. Astron. Belgrade 50, 111–114.

32. E. Simonneau, O.Atanacković –Vukmanović: Solution of line formation problem by use of iteration factors 1996, Journal of Applied Spectroscopy 63, N.5, 846–852.

33. O.Atanacković–Vukmanović: On recent advances in radiative transfer theory 1996, Publ. Astron. Obs. Belgrade No. 54, 17–18.

34. O. Atanacković–Vukmanović, E. Simonneau: Solution of the non-LTE transfer problem using the method of iteration factors 1996, in Low Temperature and General Plasmas, eds. M. Milosavljević and Z. Petrović, Nova Science Publishers, Inc.

35. O. Atanacković–Vukmanović, L. Crivellari, E. Simonneau: A forth-and-back implicit Λ- iteration 1997, The Astrophysical Journal 487, 735–746.